JG Asset Services is recommending that you invest $1,275 in a 5-year certificate of deposit (CD) that pays 3.5% interest, compounded annually. How much will you have when the CD matures? Select the correct answer. a. $1,501.10 b. $1,518.70 c. $1,505.50 d. $1,509.90 e. $1,514.30
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=1,275*(1.035)^5

=1,275*1.18768631

=$1514.30(Approx).